# An Industry Interview Study of Software Signing for Supply Chain Security

### v1

@@ Abstract
Many software products are composed by the recursive integration of components from other teams or external parties. Each additional link in a software product’s supply chain increases the risk of the injection of malicious behavior. To improve supply chain provenance, many cybersecurity frameworks, standards, and regulations recommend the use of software signing. However, recent surveys and measurement studies have found that the adoption rate and quality of software signatures are low. These findings raise questions about the practical application of software signing, the human factors influencing its adoption, and the challenges faced during its implementation. We lack in-depth industry perspectives on the challenges and practices of software signing.
To understand software signing in practice, we interviewed 18 high-ranking industry practitioners across 13 organizations. We provide possible impacts of experienced software supply chain failures, security standards, and regulations on software signing adoption. We also study the challenges that affect an effective software signing implementation. To summarize our findings: (1) We present a refined model of the software supply chain factory model highlighting practitioner’s signing practices; (2) We highlight the different challenges – Technical, Organizational, and Human — that hamper software signing implementation; (3) We report that expert subjects disagree on the importance of signing; (4) We describe how failure incidents and industry standards affect the adoption of software signing and other security techniques. Our findings contribute to the understanding of software supply chain security by highlighting the impact of human and organizational factors on Software Supply Chain risks and providing nuanced insights for effectively implementing Software Supply Chain security controls — towards Software signing in practice.
@@ 1 Introduction
Like all engineered products, software can be built more quickly when components are reused from other sources [1, 2]. This integration accelerates production but requires trusting external parties [3]. This trust increases when the provenance of components is known [4]. The strongest guarantee of provenance is a cryptographic signature created using a secret known only to the vendor. This technique, known as software signing, is underutilized by many software producers [5, 6]. However, as a result of software supply chain attacks such as SolarWinds, signing has become integral to many government, academic, and industry proposals to secure software supply chains [7, 8, 9].
Although software signing has desirable properties, it must be adopted carefully to bring benefits. At present, measurement studies have shown that missing or erroneous signatures are common across the software supply chain [5, 10, 11, 6, 12]. When signing is omitted or done unsuccessfully, researchers have demonstrated exploits [8, 13, 14, 15]. Existing security frameworks recommend software signing, but the ambiguous guidelines across these frameworks lead to a lack of a detailed implementation model [16, 17]. In order to adopt software signing, organizations would benefit from a detailed understanding of industry practices and challenges. The academic and grey literature does not offer detailed industry perspectives on the experiences that engineers have when adopting software signing.
Our work represents the first qualitative study of software signing in industry practice. We interviewed 18 industry security experts and decision-makers across 13 organizations. Our interviews focused on four aspects: the software signing practices employed in industry; the importance accorded to software signing in practice; the influence of industry standards and failure events; and the organizational, technical, and other challenges that hamper the effective implementation of software signing in practice.
We present a comprehensive summary of our findings. We refined the software supply chain factory model to show the roles of signing. While artifacts are frequently signed, signature verification is notably low. Our study reveals that the most common challenges in practice are technical, organizational, and human factors. These include key management and infrastructure, compatibility, a lack of signature verification, expertise in setting up signing tools, and developer attitudes. Our research indicates that software signing is typically used in one of three ways to mitigate risks: establishing provenance, serving as an auxiliary security technique, or fulfilling requirements akin to ticking a checkbox. While software supply chain failures, regulations, and other industry standards do influence the adoption of software signing and other security techniques, our participants have identified several issues associated with these regulations and standards.
Our contributions are:
@@
    We refined the software supply chain factory model to reflect diverse industry practices and the forms that signing may take therein (Figures 4 and III).
@@
    We identified common technical, human, and organizational challenges in adopting signing (Table IV).
@@
Significance: Software signing is a widely recommended technique, and has become mandatory for US federal software vendors. Many organizations are therefore seeking to integrate software signing into their engineering processes. This study is the first qualitative work on software signing in industry. We provide insights on software signing in industry: adoption practices, challenges, and human and organizational factors to consider. Such knowledge will guide organizations in their software signing journey, and contribute to the software engineering community’s strategies for enhancing software supply chain security.
@@ 2 Background & Related Works
Here, we cover software supply chains (section 2.1), cybersecurity considerations (section 2.2), and empirical studies (section 2.3).
@@ 2.1 Software Supply Chains
Software production is often depicted using variations on the early Software Factory Model [18, 19] or the more recent Software Supply Chain Factory Model (SSCFM) [20, 21], depicted in Figure 1. The SSCFM approximates software product from the perspective of one software engineering organization or team. The SSCFM model shows that a software product integrates first-party and third-party components, that these components are packaged for downstream use, and that the software product is at risk of malicious code injection from internal or external parties.
Figure 1: The Software Factory Model [20].
The study of software supply chains has been driven by engineering organizations’ greater reliance on external components. Many works have reported on this in various software contexts. For example, a 2024 Synopsys report stated that, in 1,067 projects across 17 industry sectors, $\sim$96integrated open-source components, and $\sim$77of total code is from open source [22]. Even in government- and safety-critical systems, the use of third-party open-source components has moved from anathema to normal [23, 24]. Analyses of open-source software package registries have shown large increases year-over-year in the number of available components, the number of actively-used components, and the number of transitive dependencies [25, 26, 27, 28].
When an external component contains a vulnerability, that vulnerability may be propagated to its dependents. The increased reliance on external components in software production has amplified the potential for exploitation of these vulnerabilities, and many attack scenarios have been proposed [29, 30, 31, 32, 33]. For example, Lauinger et al. [34] report that $\sim$37of the most popular Internet websites rely on at least one library with a publicly acknowledged vulnerability. Stock et al. showed that among 1,273 real-world client-side XSS vulnerabilities, 273 cases occurred exclusively within third-party code [35]. Upstream in the supply chain, Zimmerman et al.’s analysis of npm indicated that up to 40% of all packages depended on code with at least one publicly known vulnerability [26]. Given the prevalence of such inherited vulnerabilities, Zahan et al. [36] and similar studies have summarized attack vectors and potential weak links in software supply chains [37, 38, 39, 40, 41, 42, 43, 44].
@@ 2.2.1 Methods and Practices
Numerous methods and tools have emerged to improve software supply chain security. Okafor et al. [45] characterize them as promoting three distinct dimensions: transparency (knowing all components in the chain), separation (isolating components), and validity (ensuring integrity of components). For example, transparency is increased by adoption of Software Bills of Materials (SBOMs) [46, 47]. Separation is promoted by techniques such as sandboxing [48, 49, 50, 51, 52] and zero-trust architectures [53, 54]. Validity is improved by reproducible builds [55, 56], software signatures [5], and attestations about artifact metadata such as authorship [57].
Security standards, frameworks, and regulations have been established to complement and direct these security efforts. Notable examples include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [20], the Cloud Native Computing Foundation’s (CNCF) Software Supply Chain Best Practices [58], Microsoft’s Supply Chain Integrity Model (SCIM) [59], and the US National Institute of Standards and Technology’s (NIST) Secure Software Development Framework (NIST-SSDF) [60] and NIST 800-204D [61], which were developed in conjunction with US Executive Order 14028 (EO-14028) [62]. At present, national regulations are inadequate to force compliance with secure supply chain practices [63].
@@ 2.2.2 Software Signing in Detail
Software signing is a formally guaranteed method of establishing the authorship of a software component [64, 65, 66, 14]. Signing ensures the provenance of components through public-key cryptography (Figure 2), improving validity (provenance) in supply chains [45]. Many standards and government regulations recommend signing as a base Software Supply Chain security technique [20, 7, 67, 59, 9]. For example, in The Linux Foundation’s SLSA framework, software signing is needed for Build Security Level 3 [20].
2.3 Empirical Work on SW Supply Chain Security
The theoretical properties of techniques to improve software supply chain security are well known, whether via transparency, separation, or validity. However, adopting them in practice remains a challenge. Empirical/in vivo studies on software supply chain security have examined adoption practices and challenges to identify gaps in theoretical guarantees or usability. Some studies examine the general activities of component management, e.g., studies on dependency selection [68] and maintenance [69], and examinations of security challenges and incident handling procedures [70]. Others have examined specific techniques for supply chain security. For example, with respect to the transparency property, researchers have studied the practices of SBOM adoption [71, 72], and reported gaps in tooling and interoperability as well as incomplete adoption and maintenance across the software development lifecycle.
With respect to software signing, researchers have examined gaps in usability and adoption. Older implementations of signing were complex to use [73, 74], inspiring newer approaches with simplified tooling and support for key management [75, 76, 77]. About 25% of publicly-reported attacks from 2010-2021 exploited the missing or erroneous use of signing [15]. Gokkaya et al. showed that the top signing-related attack vectors were self-signed code, broken signature systems, and stolen certificates [42]. Several reports show that signing adoption rates for open-source components are low. Zahan et al. found that $\leq$0.5of packages in the npm and PyPI registries had signed releases [6]. Schorlemmer et al. corroborated this proportion, further noting that signing adoption is correlated with registry policies and that the correctness or quality of signatures depends on usable tooling [5]. Ladisa et al.’s survey shed light on these measurements, finding that software signing is valued but is considered moderately costly to set up [44, 78].
@@ 3 Knowledge Gaps & Research Questions
In our analysis of the literature discussed in section 2.3, we observed that our empirical understanding of software signing challenges and adoption is primarily derived from open-source software components. The available data on industry practices are surveys [44, 78]. We lack in-depth qualitative descriptions of industry experiences, creating a gap in our understanding of software supply chain security. We do not know which signing strategies are effective, the challenges associated with implementing these methods, the roles played by suggested methods and tools, and the factors influencing their adoption. To fill this gap, we ask four research questions:
@@
    How and why is software signing implemented by software teams?
@@
    How does software supply chain failures and standards influence the adoption of software signing?
@@ 4 Methodology
We conducted semi-structured interviews with subjects who were cybersecurity practitioners. We depict our methodology in Figure 3. In this section we discuss our research design rationale (section 4.1), protocol development (section 4.2), participant recruitment (section 4.3, data analysis (section 4.4), and study limitations (section 4.5).
@@
4.1 Approach — Research Design Rationale
Our questions require insight into practitioner perspectives. We considered two factors in selecting a methodology. First, since the prior work in this space is limited, there was relatively little basis on which to structure a closed-ended survey. Second, we expected recruitment to be a problem, since the target population has high security expertise and a broad organizational view, and thus are likely to be high-ranking (and busy) people. We therefore opted for a semi-structured interview [79], which would permit us to capture emergent results [80] (factor 1), and to leverage our professional relationships to improve response rates (factor 2).
The resulting data is expected to be exploratory, and rooted in practitioners’ perspectives and experiences [79]. The phenomenon of interest is software signing as a security method, which is a contextual decision based on an organization’s distinct policies, security expectations, perceived threats, and customer requirements [81]. We therefore adopt an interpretivist perspective for our research methodology design and analysis. An interpretivist approach treats each situation as unique [82], focusing on in-depth variables and factors related to each context [83].
@@ 4.2 Protocol Design and Development
To create the initial interview protocol, we studied three kinds of resources. First, to ensure a common language and a comprehensive set of topics, we studied software supply chain security frameworks such as SLSA, CNCF, SCIM, and NIST-SSDF (section 2.2). These frameworks use variants of the Software Supply Chain Factory Model (SSCFM, cf. Figure 1). We therefore used the SSCFM as our reference framework for structuring the C Sections of our interview protocol. This choice allowed us to investigate signing practices at various stages of the SDLC while using a model familiar to our subjects. Second, we examined grey literature sources to gain insights into some practical uses of software signing. Our primary reference was the official Sigstore blog as of Jan. 31, 2023. This resource has eight user reports about adopting software signing in various industry contexts, from which we derived questions about the contextual and organizational factors that subjects consider. Third, we reviewed adjacent literature [68, 69, 84] to understand effective questions related to software supply chain security, while avoiding duplication.
@@
After protocol development, we refined the protocol through practice interviews (2) and pilot interviews (2). For practice, the lead author interviewed two of the secondary authors. These practice sessions led to modifications in sections B and C of our interview protocol (see Appendix Appendix C). After recruiting subjects, we treated the first two subjects as pilot interviews. During these pilots, we gathered additional insights into how subjects perceived software supply chain attacks and reorganized the protocol, particularly focusing on questions related to the perceived sources of greatest risk (section B). Since few questions were changed or added, we included the data from the pilot interviews in our study where feasible. The final change occurred after the seventh interview. The author team discussed the results so far and identified one emergent topic: how practitioners’ organizations integrate feedback into their SSC security processes. One question was added for use in the remaining interviews.
Our interview protocol is summarized in Table I, with a full version in Appendix Appendix B.
@@
TABLE I: Summary of interview protocol. See Appendix Appendix B for full protocol. Topic D was not analyzed in this study.

### v2

@@ Abstract
Many software products are composed of components integrated from other teams or external parties. Each additional link in a software product’s supply chain increases the risk of the injection of malicious behavior. To improve supply chain provenance, many cybersecurity frameworks, standards, and regulations recommend the use of software signing. However, recent surveys and measurement studies have found that the adoption rate and quality of software signatures are low. We lack in-depth industry perspectives on the challenges and practices of software signing.
To understand software signing in practice, we interviewed 18 experienced security practitioners across 13 organizations. We study the challenges that affect the effective implementation of software signing in practice. We also provide possible impacts of experienced software supply chain failures, security standards, and regulations on software signing adoption. To summarize our findings: (1) We present a refined model of the software supply chain factory model highlighting practitioner’s signing practices; (2) We highlight the different challenges–technical, organizational, and human–that hamper software signing implementation; (3) We report that experts disagree on the importance of signing; and (4) We describe how internal and external events affect the adoption of software signing. Our work describes the considerations for adopting software signing as one aspect of the broader goal of improved software supply chain security.
@@ 1 Introduction
Like all engineered products, software is built more quickly when components are reused from other sources [75, 40]. Although this integration accelerates production, it requires trusting external parties [5]. This trust increases when the provenance of components is known [68], such as via software signing. Although historically underutilized [64, 97], due to attacks such as SolarWinds [31], signing has become part of many government, academic, and industry proposals to secure software supply chains [80, 19, 17].
Although software signing has desirable effects, it must be adopted carefully to bring benefits. At present, measurement studies show that missing or erroneous signatures are common across the software supply chain [64, 53, 42, 97, 93]. Existing security frameworks recommend software signing, but the general guidelines in these frameworks omit a detailed implementation model [61, 100]. In order to adopt software signing, organizations would benefit from a detailed understanding of industry practices and challenges. This view is currently lacking in the academic and grey literature.
Our work represents the first interview study that describes software signing practices and rationales in industry. This method gives a deep look at a topic, exposing many directions for further study and implications for practice. In this work, We interviewed 18 experienced security practitioners across 13 organizations. Our interviews focused on four aspects: the software signing practices employed in industry; the challenges that hamper the effective implementation of software signing in practice; the importance accorded to software signing in practice; and the influence of failure events and industry standards.
Our results provide a detailed understanding of industry practices and challenges in using software signing to promote software supply chain security. By integrating our subjects’ descriptions, we introduce a refinement of the Software Supply Chain Factory Model that indicates the possible and actual roles of software signing. Our results reveal three kinds of challenges in adopting signing: technical, organizational, and human aspects. Additionally, our findings indicate that practitioners hold varying views on the importance of signing for supply chain security. Lastly, we found that security events — both major failures such as SolarWinds, and new standards and regulations — appear to have a minimal effect on the adoption of software signing.
In summary, we make the following contributions:
@@
    We refined the software supply chain factory model to reflect diverse industry practices and the forms that signing may take therein (section 5.1).
@@
    We identified common technical, human, and organizational challenges in adopting signing (section 5.2).
@@ 2 Background & Related Works
Here, we cover software supply chains (section 2.1), cybersecurity considerations (section 2.2), and relevant empirical studies (section 2.3).
@@ 2.1 Software Supply Chains
Software production is often depicted using variations on the early Software Factory Model [47, 14] or the more recent Software Supply Chain Factory Model (SSCFM) [79, 1], depicted in Figure 1. The SSCFM approximates software product from the perspective of one software engineering organization or team The SSCFM model illustrates that a software product integrates both first-party and third-party components, which are subsequently packaged for downstream use. These integration points also represent potential vulnerabilities where the software product is at risk of malicious code injection from either internal or external sources.
Figure 1: The Software Supply Chain Factory Model [79].
The study of software supply chains has been driven by engineering organizations’ greater reliance on external components. Many works have reported on this reliance in various software contexts. For example, a 2024 Synopsys report stated that, in 1,067 projects across 17 industry sectors, $\sim$96integrated open-source components, and $\sim$77of total code is from open source [76]. Even in government and safety-critical systems, the use of third-party open-source components has moved from anathema to normal [32, 22]. Analyses of open-source software package registries have shown large increases year-over-year in the number of available components, the number of actively-used components, and the number of transitive dependencies [91, 99, 38, 54].
The increasing reliance on external components in software production amplifies the potential for the exploitation of vulnerabilities that may be contained in these components. As an example of this risk, Lauinger et al. [46] report that $\sim$37of the most popular Internet websites rely on at least one library with a publicly acknowledged vulnerability. Given the prevalence of inherited vulnerabilities, Zahan et al. [98] and similar studies have summarized attack vectors and potential weak links in software supply chains [25, 9, 71, 96, 52, 29, 56, 43].
@@ 2.2.1 Methods and Practices
Numerous methods and tools have emerged to improve software supply chain security. Okafor et al. [57] state that these approaches promote three distinct dimensions: transparency (knowing all components in the chain), separation (isolating components), and validity (ensuring integrity of components). For example, transparency is increased by adoption of Software Bills of Materials (SBOMs) [51, 95]. Separation is aided by techniques such as sandboxing [74, 4, 45, 83, 82] and zero-trust architectures [73, 8]. Validity is improved by reproducible builds [44, 86], software signatures [64], and attestations about artifact metadata (e.g., authorship) [81].
Security standards, frameworks, and regulations have been established to complement and direct these security efforts. Notable examples include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [79], the Cloud Native Computing Foundation’s (CNCF) Software Supply Chain Best Practices [2], Microsoft’s Supply Chain Integrity Model (SCIM) [3], and the US National Institute of Standards and Technology’s (NIST) Secure Software Development Framework (NIST-SSDF) [18]. Current national regulations do not force compliance with secure practices [48].
@@ 2.2.2 Software Signing in Detail
Software signing is a formally guaranteed method of establishing the authorship of a software component [28, 27, 66, 20]. It ensures the provenance and integrity of components in the supply chain through public-key cryptography. The process involves creating a hash of the software, encrypting it with the author’s private key, and attaching the resulting signature to the software. When users download the software, they decrypt the signature with the corresponding public key and compare the hash to ensure the software has not been altered. A typical workflow for a package signing is shown in Appendix A.
Keyless signing[53, 35] eliminates the need for authors to manage long-term private keys. Instead, authors authenticate with an identity provider, and an ephemeral key associated with their verified identity is issued for each signing. Users then additionally verify that the public key used for the software verification is associated with the intended author, ensuring both software integrity and authenticity.
Many standards and regulations recommend signing as a base Software Supply Chain security technique [79, 80, 70, 3, 17]. For example, in The Linux Foundation’s SLSA framework, software signing is part of Build Security Level 3 [79].
2.3 Empirical SW Supply Chain Security
Although many theoretical techniques for software supply chain security are known, adopting them in practice remains a challenge. Empirical/in vivo studies on software supply chain security have examined adoption practices and challenges to identify gaps in theoretical guarantees or usability[90, 62, 29]. Some studies examine the general activities of component management, e.g., dependency selection [60] and maintenance [11], and examinations of security challenges and incident handling procedures [88]. Others have examined techniques for supply chain security. For example, for the transparency property  section 2.2.1, researchers have studied the practices of SBOM adoption [94, 41] and reported gaps in tooling, interoperability, adoption, and maintenance. For separation, studies have examined the performance of sandboxing [87] and its adoption in open-source ecosystems [7]. For validity, reproducible builds have been analyzed in PyPi [86].
Concerning software signing, researchers have examined challenges in its usability and adoption that create opportunities for attacks [90, 62]. In 2010–2021, roughly 25% of publicly reported attacks exploited the missing or erroneous use of signing  [36]. Top signing-related attack vectors include self-signed code, broken signature systems, and stolen certificates  [29]. Although software signing is valued, it is considered costly to set up, as reported in Ladisa et al.’s survey  [43, 89]. Several reports show that signing adoption rates for open-source components are low. For instance, Zahan et al. found that $\leq$0.5of packages in the npm and PyPI registries had signed releases  [97]. Schorlemmer et al. corroborated this proportion, further noting that signing adoption is correlated with registry policies and that the correctness or quality of signatures depends on usable tooling  [64].
@@ 3 Knowledge Gaps & Research Questions
In our analysis of the literature discussed in section 2.3, we observed that our empirical understanding of software signing challenges and adoption is primarily derived from open-source software components. The available data on industry practices are surveys [43, 89]. We lack in-depth qualitative descriptions of industry experiences, creating a gap in our understanding of software supply chain security. We do not know which signing strategies are effective, the challenges associated with implementing these methods, the roles played by suggested methods and tools, and the factors influencing their adoption.
To fill this gap, we ask four research questions:
@@
    Where and how is software signing implemented by software teams?
@@
    How do internal and external signing events influence the adoption of software signing?
@@ 4 Methodology
We conducted semi-structured interviews with subjects who were cybersecurity practitioners. We depict our methodology in Figure 2. We describe our research design (section 4.1), protocol development (section 4.2), participant recruitment (section 4.3), data analysis (section 4.4), and limitations (section 4.5). Our Institutional Review Board (IRB) approved this study.
@@
4.1 Research Design and Rationale
Our research questions require insight into practitioner perspectives. We considered two factors in selecting a methodology. First, because the prior work in this space is limited, there was relatively little basis on which to structure a closed-ended survey. Second, our research questions warrant long-form descriptive-type answers that have to be extracted through open-ended questions, opinions, and detailed observations. We therefore opted for a semi-structured interview [63], which would permit us to capture emergent results [21].
The resulting data is expected to be exploratory, and rooted in practitioners’ perspectives and experiences [63]. The phenomenon of interest is software signing as a security method, which is a contextual decision based on an organization’s distinct policies, security expectations, perceived threats, and customer requirements [69]. We therefore adopt an interpretivist perspective for our research methodology design and analysis. This approach treats each situation as unique [10], focusing on in-depth variables and contextual factors [6]. Our choice of the interpretivist framework was also motivated by our sampling technique as we discuss later (section 4.3).
@@ 4.2 Protocol Design and Development
We studied three kinds of resources as we created the initial interview protocol.
  1. 1.
@@
    Software Supply Chain Security Frameworks: To ensure a common language and a comprehensive set of topics, we studied software supply chain security frameworks such as SLSA, CNCF, SCIM, and NIST-SSDF (section 2.2). These frameworks use variants of the Software Supply Chain Factory Model (SSCFM, cf. Figure 1). We therefore used the SSCFM as our reference framework for structuring the C Sections of our interview protocol. This choice allowed us to investigate signing practices at various stages of the SDLC while using a model familiar to our subjects.
  2. 2.
    Grey Literature: To gain insights into practical uses of software signing, we examined grey literature such as industry blogs [30, 67], whitepapers [16], and case studies [23, 58, 93]. These sources documented the adoption of software signing in various industry contexts, and, from these, we derived questions about the contextual and organizational factors that subjects consider. We collected these grey literature sources through Web searches of software signing tools mentioned or recommended by any of the security frameworks we studied. Additionally, we reviewed linked articles related to software signing in these standards.
  3. 3.
    Background & Literature Review: Our background and review of related works (section 2) provided a foundation for knowing what questions would be effective for answering our research questions while avoiding duplication. Broadly, our literature review process, which yielded all works cited in section 2) was based on a forward and backward snowball search. We began with works published in prominent cybersecurity and software engineering venues (USENIX, CCS, IEEE S&P; ICSE, FSE, ASE) and covered a 10-year period.
@@
After protocol development, we refined the protocol through practice interviews (2) and pilot interviews (2). For practice, the lead author interviewed two of the secondary authors. These interviews led to modifications in sections B and C of our interview protocol (see Appendix F). After recruiting subjects, we treated the first two subjects as pilot interviews. During these pilots, we gathered additional insights into how subjects perceived software supply chain attacks and reorganized the protocol, particularly focusing on questions related to the perceived sources of greatest risk (section B). Since few questions were changed or added, we included the data from the pilot interviews in our study where feasible  [^1] We included data from the pilot interviews in our analysis, as qualitative methods guidance supports this practice when the protocol demonstrates its ability to effectively answer research questions  [78, 37]. This approach is particularly justifiable in contexts where recruitment is challenging, as was the case here and in similar work  [49]. Quantitatively, we observed that: (1) the pilot interviews used nearly all ( 92%) of the final protocol, as only 3 out of 35 questions were changed afterward, and (2) the unique codes derived from the pilot interview transcripts were comparable in quantity and richness to those from other participants (see  Figure 5).. The final change was made after the seventh interview. The author team reviewed the results up to that point and identified an emergent topic: how practitioners’ organizations integrate feedback into their software supply chain (SSC) security processes. Consequently, one additional question was included in the remaining interviews, to address this emerging topic.
Defining Internal & External Events to Answer RQ4: Unlike the other research questions, RQ4 focuses on investigating how internal and external factors influence organizational decisions to adopt software signing. Following the definitions of Schorlemmer et al. [64], we scope this part of the protocol to the impact of (1) Software Supply Chain failures, and (2) Software Supply Chain standards.
Table 1 summarizes our interview protocol. The full protocol is given in Appendix E.
@@
Table 1: Summary of interview protocol. Topic D is not analyzed in this paper. See Appendix E for full protocol.